OpenStudy (scarlettthecamel):

Solve the simultaneous equations : xy=3, y=x-2

rishavraj (rishavraj):

plug y = x - 2 in the equation xy = 3

OpenStudy (scarlettthecamel):

x^2-2=3

rishavraj (rishavraj):

hmmmm nah it would be x(x - 2) = 3 so x^2 - 2x = 3

OpenStudy (scarlettthecamel):

ohh right

OpenStudy (scarlettthecamel):

i don't understand where this is leading to

OpenStudy (scarlettthecamel):

i need to find 2 x's and 2 y's

rishavraj (rishavraj):

on solving x^2 - 2x - 3 = 0 u will get two values of "x" plug them in equations and get corresponding "y"

OpenStudy (scarlettthecamel):

okay ill try plugging x in

OpenStudy (scarlettthecamel):

i really don't know there's only one value of x

rishavraj (rishavraj):

on solving x^2 - 2x - 3 = 0 -_- u will get x = -1 , 3 now plug value of x in any of the two equation and get corresponding "y"

OpenStudy (scarlettthecamel):

-2-9-2x-3=0

rishavraj (rishavraj):

@scarlettthecamel HMMMM ???? on plugging x = -1 y = x - 2 y = -1 -2 y = -3

rishavraj (rishavraj):

similarly plug x = 3

OpenStudy (scarlettthecamel):

i really don't get any of this

rishavraj (rishavraj):

see first of all its given y = x - 2 so plugging y in xy = 3 u get x(x - 2) = 3 solve the equation for "x"